OpenStudy (anonymous):

sin 3pi/8 use half angle formula to find exact value. please show work

OpenStudy (callisto):

\[cos(2\theta) = 1-2sin^2\theta\] So, \[cos(2 \times\frac{3\pi}{8}) = 1-2sin^2\frac{3\pi}{8}\]Left side is easy to simplify (special angle), then you can find sin(3pi/8)